OpenStudy (anonymous):

The expression \sqrt[5]{v^{25}}

OpenStudy (anonymous):

Does it ask to simplify the expression?

OpenStudy (anonymous):

yes i need to simplify the expression

OpenStudy (anonymous):

\[ v^{25} = (v^5)^5 \]

OpenStudy (anonymous):

\[ v^{25} =v^{5\times 5} = (v^5)^5 \]So:\[ \sqrt[5]{v^{25}} = \sqrt[5]{(v^5)^5} = v^5 \]

OpenStudy (anonymous):

In general: \[ \sqrt[a]{v^b} = v^{b/a} \]Which in this case where \(a=5\) and \(b=25\): \[ \sqrt[5]{v^{25}} = v^{25/5} = v^5 \]

OpenStudy (anonymous):

thanx